After training, Mary Fernandez, a computer technician, had an average observed time for memory-chip tests of 12 seconds. Mary’s performance rating is 100%. The firm has a personal fatigue and delay allowance of 15%.

A) Find the normal time for this process

B) Find the standard time for this process
Answer #1

A)

Observed time \(=12\) seconds Performance rating factor \(=100 \%\) \(=100 / 100\)

\(=1\)

Normal time \(=\) Average observed time \(\times\) Factor of performance rating \(=12 \times 1\)

\(=12\) seconds

B)

\(\begin{aligned}&\text { Allowance factor }=15 \% \\&\qquad \begin{aligned} &=15 / 100 \\ &=0.15 \end{aligned} \\&\text { Standard time } & =\frac{\text { Normal time }}{1-\text { Allowance factor }} \\&\text { Standard time } & =\frac{12}{1-0.15} \\& & =\frac{12}{0.85} \\& & =14.117 \text { seconds }\end{aligned}\)